What is the density of helium gas at 25 degree celsius and 760 torr?

R= 0.082 Lxatm/molxk
AW= 4 g/mole

a) 0.16 g/L

b) 6.1 g/L

c) 124 g/L

d) 0.082 g/L
1
Expert's answer
2014-06-18T09:04:25-0400
According to the ideal gas law,
PV = nRT
n = m/M - number of moles
d = m/V - density
Hence,
P = (m/MV)RT
PM = dRT
d = PM/RT
760 torr = 1 atm
d = (1 atm x 4 g/mol)/(0.082 Lxatm/molxK x 298 K) = 0.164 g/L
So, the correct option is (a): 0.16 g/L